# On Creativity and Open-Endedness

### v1

@@ Introduction
Many process theorists consider transformational creativity as the most profound [cf. Lamb et al., 2018], including Boden herself, who notes that only transformational creativity allows for the most radical forms of surprise. She considers surprise to be a defining feature of creativity, next to novelty and value.
Taylor notes that, since Banzhaf et al.’s [2016] types are defined in terms of whether or not the novelty results in a change in the model or meta-model of a system, it closely resembles the three different kinds of creativity as outlined by [Boden, 2004]. Exploratory novelties result from applying the current dynamics of a system (summarised as a model) to multiple possible states. Expansive novelties result from applying similar models that share concepts present in the current meta-model. Lastly, transformational novelties can only result from models made of entirely new concepts, which necessitates a change in the meta-model description of the system’s dynamics. Taylor directly translates these three types of novelties into three types of open-ended processes that produce these novelties. In particular, the difference between I-creativity (previously P-creativity, when an idea, action, or object is new to an individual [Boden, 2015]) and H-creativity (when an idea, action, or object is new to historically [Boden, 2004]) is compared to the difference between transformational and expansive novelties, respectively.
Taylor [2021] argues that systems that produce expansive and transformational novelties and innovations can be generalized further into transdomain bridges and non-additive compositional systems [Taylor, 2021]. By characterizing a dynamical system by its parameters, spatial arrangement, and dynamics (Parameter Space–Organisation Space–Action Space (POA)), all evolutionary innovations, including the three mentioned in [Taylor, 2019], are either a physical phenomenon that had previously been unseen, or resulting from new, persistent systems constructed from within the environment. This last point is related to the concept from creativity research that involves transforming the space in which innovations can be explored. By explicitly considering an environment that can host new physical systems– and therefore new dynamics and new spaces of possibility– the resulting non-additive compositional systems make it possible for evolutionary systems to expand their search spaces. By doing so, Taylor suggests that evolutionary systems can evolve in ways that are expansive and transformational, not just exploratory.
@@
Song [2022] recites these previous connections and highlights a “loose [conceptual] equivalence” between creativity (as defined by Boden [2004] in terms of surprise, novelty, and usefulness) and OE (as defined in terms of novelty, diversity, and complexity). Song grounds this equivalence in Schmidhuber’s (2008) theory of beauty-as-compressibility. Compressibility in turn is often used to estimate complexity in OE [e.g. Earle et al., 2021]. Notably, Schmidhuber’s theory has received arguably little approval, and Song’s attempt at an equivalence misses instances of creativity that are disconnected from beauty. This leaves the exact relationship between compressibility and these three components of creativity open. Song also notes that requirements for OE are weaker than for creativity according to Boden’s definition, in that OE “requires one of novelty, diversity, and complexity, while having the contingency for the other qualities”. Our goal is to go beyond statements of “loose equivalence” and illuminate the exact nature of this match, based on the wide array of different conceptualisations of OE and creativity.
@@
To this end, as our second contribution, we put forward a preliminary list of questions, each of which proposes a potential connection between OE and creativity. We invite researchers to dig into these proposals to identify similarities, differences and potential extensions to the concepts and measurement of OE and creativity. These questions were identified by the authors, who are from the ALife (initials anonymized) and CC (initials anonymized) communities, respectively, of which at least two (initials anonymized) have actively contributed to research in both fields. In eight remote sessions, we collaboratively constructed an affinity diagram, capturing the connections between the two concepts. Affinity diagrams represent a qualitative data analysis method used to group data into emergent categories [Sharp et al., 2019]. Akin to self-organized focus groups [Morgan, 1996], we proposed, explored, and discussed potential connections between OE and creativity along several dimensions, such as agency, timescale, substrate, and subjectivity.
@@
This process was informed by the authors’ detailed knowledge of each field’s research heritage. As a foundation for our investigation, we first reviewed existing surveys of OE and (computational) creativity, as well as existing work on their connections, as discussed throughout this paper. To summarize these connections, we translate each into an open research question for future research directions. We refer to existing work to detail and refine these questions:
@@
  • Could generative creativity [Bown, 2012] or the Pagan creativity tradition [Still and d’Inverno, 2016] provide the most universal similarity to OE since it requires novelty (but not value), and its inspiration from natural processes?
@@
  • Could the notion of adaptive creativity [Bown, 2012; Guckelsberger et al., 2017] provide a closer fit to cases of OE that comprise a notion of value as adaptive success?
  • Value is mostly considered an essential component of creativity [Runco and Jaeger, 2012]; are any conceptualisations of value exclusive to creativity or OE and thus deny a strict equivalence match between both concepts?
@@
  • Do the modeling approaches of dynamic processes in CC and OE differ in their requirement for continuity? More specifically, would CC researchers consider a system to be creative even if it quickly exhausted its potential of generating novel and valuable artifacts? To this end, what does the scarce work on continuous CC [Cook, 2021] betray?
@@
  • Could Boden’s [2004] three types of creative processes and Taylor’s [2021] ALife interpretation be used as diagnostics for OE? Would a system be considered open-ended even if it only briefly exhibited transformational-, but then mostly produced explorative novelties?
@@
  • Is creative autonomy [Jennings, 2010; Saunders, 2012; Guckelsberger et al., 2017] a necessary requirement for closing the gap between creativity and OE?
@@
  • Can research on OE and CC inspire the viewpoint from which the respective phenomenon is driven and evaluated? Most CC systems are evaluated w.r.t. the goals of an external observer (potentially injected into the system), and creativity is driven from the inside-out, rather than through the system’s surroundings.
  • What is the trade-off between OE and CC, and controllability in a dynamical system? How are these two concepts related to an observer’s ability to control a system’s dynamics? Can the relationship of either one of the first two concepts illuminate the role of the other?
@@
The dual pursuits of synthesizing creativity in CC and OE in ALife involve understanding the rich histories of each research agenda, as briefly summarized in this short paper. Our goal has not been to move toward a complete coalescence of creativity and OE but, through intentional dialogue about both the similarities and differences in approaching each challenge, we hope to arrive at a clearer vision of how best to further our individual aims by understanding as tightly interconnected with those of CC and ALife as neighbouring communities. To this end, we suggest a few questions that, if attended, would likely help push both communities toward a fuller understanding of these respective fields:
  • Can we translate research facets and goals from CC to OE research, and vice versa? For instance, an interesting exercise might be to imagine “creativity as it might be”.
@@
A long history of exchange between ALife and CC [Saunders, 2012; Bown, 2012; McCormack, 2012; Guckelsberger et al., 2017; McCormack, 2019] bears witness of this endeavour’s inspirational and transformational potential.
@@
The main goal of this paper is to revisit and collect several connections between CC (CC) and OE (OE) from the literature and to recommend new avenues of interdisciplinary directions for future research in creativity research, AI (AI), and ALife (ALife). Since CC and OE have been identified as the “final frontier” and “the last grand challenge” of AI research, respectively, we stress that there is opportunity for both communities to learn from each other’s research heritage, progress, and challenges.
To this end, we provide a short, comprehensive, and historical context for progress in both research areas. Special attention is placed on existing literature that connects concepts from creativity and OE explicitly. We discuss potentially fertile questions for both fields to consider, such as whether there exist any nontrivial differences between creativity and OE or whether they are fundamentally the same concept. We also identify possible synergy effects of a cross-community research agenda that encompasses both, including potential strategies to reduce ambiguity in defining and measuring each concept.
While this provides context on much of the progress made on the intersection of both areas, we most importantly sound the call for further discussion and collaboration between the (computational) creativity and ALife open-endedness communities. It is crucial to establish a dialogue between them, since doing so might provide additional context, insight, and directions moving forward.
While we do not intend to resolve these connections conclusively within this limited space and team of authors, we provide a starting point for dedicated interdisciplinary investigations. We hope that such endeavors will deepen our understanding of both research challenges and further elucidate the requirements to reproduce them in artificial systems– including reducing open conceptual ambiguities and fostering collaborative research between these two communities.
@@ Acknowledgements
We thank our anonymous reviewers for their excellent feedback and support for this paper’s mission.

### v2

@@ Introduction
Boden [2015a] emphasizes that these “are analytical distinctions between different psychological/biological processes. They are not intended as overall descriptions of the resulting forms” (or product, or concept, or artifact) resulting from a process, as a particular form may have been created by multiple types of processes. Many process theorists consider transformational creativity as the most profound [cf. Lamb et al., 2018], including Boden herself, who notes that only transformational creativity allows for the most radical forms of surprise. She considers surprise to be a defining feature of creativity, next to novelty and value [Boden, 2004].
Applied to ALife, combinational creative mechanisms produce new combinations of existing forms, such as new kinds of macromolecules (sugars, proteins, etc). Exploratory creative mechanisms produce variants of particular structure types that have already been expressed, such as a variations of a type of protein. In other words, it identifies categories or classes of objects and explores variants in that particular category. Finally, transformational creative mechanisms produce forms that are impossible under a pre-supposed possibility space but are physically realized under a new transformational possibility space. Especially the last type is of interest to ALife researchers because it is closely related to discussions about OE processes that continually transform the adjacent possible space of forms that are physically realizable [Kauffman, 2019].
While Boden’s [2015a] account considers ALife more generally, Soros and Stanley [2016] soon thereafter drew more specific connections to OEE (but not OE). They highlight a potential connection between Taylor et al.’s [2016] primary categories of hallmarks of OEE (ongoing adaptive novelty, and ongoing growth of complexity) and Boden’s [2004] concepts of exploratory and transformational creativity. Additionally, they raise three open questions about the relationship between OEE and creativity theory: (1) (mentioned earlier) Are there nontrivial differences between creativity and OEE or are they fundamentally the same concept? (2) Is there any utility in formally pursuing the nature of the relationship between evolution and creativity? (3) Do there exist metrics and practices for studying computational creativity (or creativity in general) that could help identify and quantify hallmarks of open-ended evolution? They thus appeal to the CC and ALife struggles to define and measure creativity and OE, respectively.
Closely related, Taylor [2019] relabels and connects different kinds of novelties defined by Banzhaf et al. [2016] with Boden’s types of creativity. Banzhaf et al. distinguish three types of novelty based on whether they require changes to a system’s model or meta-model. The model describes the system’s behavior through concepts which are themselves defined in a meta-model. Exploratory novelties (Banzhaf et al.: variation) result from applying the current dynamics of a system (summarised as a model) to multiple possible states. Expansive novelties (Banzhaf et al.: innovation) result from applying similar models that share concepts present in the current meta-model. Lastly, transformational novelties (Banzhaf et al.: emergence) can only result from models made of entirely new concepts, which necessitates a change in the meta-model description of the system’s dynamics. In a footnote, Taylor [2019] notes that these three types “fit closely with Boden’s concepts of exploratory, combinational and transformational creativity”. He does not elaborate on this further, except for saying that the term compositional is preferred over Boden’s combinational “to emphasise that the size of structures may increase, and that the specific arrangement and connections between components might be important”. Crucially, Taylor interprets the three types of novelties as flavors of ongoing adaptive novelty as a hallmark of OEE [Taylor et al., 2016, also mentioned above] and consequently considers them three types of open-ended processes. He thus draws parallels between Boden’s theory and ongoing adaptive novelty only. We note that this marks a distinction to Soros and Stanley’s [2016] mapping mentioned above which also includes ongoing adaptive complexity. This ambiguity has so far not been resolved.
Also in a footnote, Taylor [2019] draws another parallel to an earlier distinction by Boden [2015a] between I- and H-creativity. I-creativity [originally P-creativity, Boden, 2004] captures the case when an idea, action, or object is new to an individual [Boden, 2015a]), while H-creativity only holds when it is historically new, to any observer. Taylor [2019] notes that Banzhaf et al. [2016] originally defined transformational novelty (Banzhaf et al.: emergence) relative to the current model and meta-model, ruling out that e.g. a major transition, which would have to be added to the meta-model, could be considered transformational again after it initially occurred. Instead, he suggests defining novelty as “relative to the initial model and meta-model applied to an evolutionary system at its inception”, and connects this change to Boden’s [2015a] distinction between I- and H-creativity.
@@
Song [2022] recites some of these previous connections in less detail and highlights a “loose [conceptual] equivalence” between creativity, as defined by Boden [2004] in terms of surprise, novelty, and value (e.g. usefulness, beauty, etc., and OE – as defined in terms of novelty, diversity, and complexity (see Background). Song grounds this equivalence in Schmidhuber’s (2008) theory of beauty-as-compressibility. Compressibility in turn is often used to estimate complexity in OE [e.g. Earle et al., 2021]. Notably, Schmidhuber’s theory has received arguably little approval, and Song’s attempt at an equivalence misses instances of creativity that are disconnected from beauty. This leaves the exact relationship between compressibility and these three components of creativity open. Despite the contentiousness of OE definitions, Song also notes that requirements for OE are weaker than for creativity according to Boden’s definition, in that OE “requires one of novelty, diversity, and complexity, while having the contingency for the other qualities” [Song, 2022].
Our goal is to go beyond statements of “loose equivalence” and propose candidates for connections which further illuminate the exact nature of this match, respecting the wide array of different conceptualisations of OE and creativity.
@@
To this end, as our second contribution, we put forward a preliminary list of questions, each of which proposes a potential connection between OE and creativity. We invite researchers to engage in in-depth exploration of these similarities, differences and potential extensions to both concepts in future work. The questions were identified by the authors, who self-identify with the ALife (LBS, AMA, OF) and CC (LBS, SK, CG) communities, and of which at least two (LBS, CG) have contributed research to both fields. In eight remote sessions, we collaboratively constructed an affinity diagram, capturing the connections between the two concepts. Affinity diagrams represent a qualitative data analysis method used to group data into emergent categories [Sharp et al., 2019]. Akin to self-organized focus groups [Morgan, 1996], we proposed, explored, and discussed potential connections between OE and creativity along several dimensions, such as agency, timescale, substrate, and subjectivity.
@@
This process was informed by the authors’ detailed knowledge of each field’s research heritage. As a foundation for our investigation, we first reviewed existing surveys of OE and (computational) creativity, as well as existing work on their connections, as discussed in the previous sections. To summarize these connections, we translate each into an open research question for future research. We reference potentially related work as a starting point where applicable:
@@
  • Could the concept of generative creativity [Bown, 2012] from CC and the Pagan creativity tradition [Still and d’Inverno, 2016] provide the closest equivalence match to OE since both conceptualisations require novelty but not value, and are inspired from natural processes?
@@
  • Could the CC concept of adaptive creativity [Bown, 2012; Guckelsberger et al., 2017] match with OE formulations that comprise a notion of value as adaptive success?
  • Value is commonly considered an essential component of creativity [Runco and Jaeger, 2012], but its interpretation is ambiguous and its context-/domain-dependency/agnosticity subject to ongoing debate in CC and creativity studies [Loughran and O’Neill, 2018]; are any specific conceptualisations of value exclusive to creativity or OE, thus denying a strict equivalence match?
@@
  • Do the goals and modelling approaches in CC and OE differ in their requirement for continuity? More specifically, would (computational) creativity researchers consider a system to be creative even if it quickly exhausted its potential of generating novel and valuable artifacts? Could ALife’s understanding of OE prompt an extension of CC’s scope toward continuously creative system? What does the scarce work on continuous CC [Cook, 2021] betray?
  • Could Wiggins[2006b; 2006a; 2019] and Ritchie’s [2012] formalisations and extensions of Boden’s [2004] theory of creative processes reduce ambiguity in mapping proposals to different types of novelties [Soros and Stanley, 2016; Taylor, 2019, highlighted earlier] as hallmark of OEE and consequently support the evaluation of OE?
  • Moreover, could their mechanistic modelling of creative processes, e.g. transformational creativity as search on a meta-level (cf. Banzhaf et al. [2016]: meta-model) inform the development of OE algorithms, e.g. those capable of emergence [Banzhaf et al., 2016]? Could this be further inspired and supported by existing and new, collaborative CC algorithm proposals to realise transformational creativity [e.g. Grace and Maher, 2015; Demke et al., 2023]?
@@
  • Could Boden’s [2004] three types of creative processes and (ideally unambiguous) ALife interpretations [Soros and Stanley, 2016; Taylor, 2019] be used as diagnostics for OE? Would a system be considered open-ended even if it only briefly exhibited transformational-, but then mostly produced exploratory novelties?
  • Is Banzhaf et al.’s [2016, distinguishing model and meta-models] or Boden/Wiggins/Ritchie’s [2004; 2006a; 2006b; 2019; 2012, distinguishing object- and meta-level] theory more expressive? Given that present formalisations of the latter only permit for certain modifications on the meta-level (e.g. to the traversal function), could Banzhaf et al.’s [2016] theory inspire extensions of the latter, or the latter inform constraints on the former?
  • Having resolved questions on conceptual (dis-)similarity, under which circumstances, if at all, can creativity/OE be used to characterize an open-ended/creative process [e.g. Bedau, 1998; Packard et al., 2019; Hughes et al., 2024]?
@@
  • Is creative autonomy [Jennings, 2010; Saunders, 2012; Guckelsberger et al., 2017] a necessary requirement for futher narrowing the gap between creativity and OE?
@@
  • CC researchers have proposed autopoiesis and adaptivity [Saunders, 2012; Guckelsberger et al., 2017] as requirements and drivers of “genuine” (not designer-imposed) creative autonomy and intentional agency. Could CC researchers adopt synthetic equivalents of these processes from ALife to progress on this agenda? Could OE research inform CC of other requirements for and drivers of creative autonomy, e.g. through the lens of continuity?
  • Can research on OE and CC inspire the viewpoint from which the respective phenomenon is driven and evaluated? Despite few counter-proposals [e.g. Guckelsberger et al., 2017], most CC systems are evaluated w.r.t. the goals of an external observer (often injected into the system) [Colton, 2008; Colton et al., 2012]. Moreover, despite the popularity of systems theories of creativity [Vygotsky, 1971; Csikszentmihalyi, 2014], creativity in most, albeit notably not all [e.g. Saunders and Bown, 2015], systems is driven from the inside-out, and decoupled from a system’s surroundings [Guckelsberger et al., 2021].
  • Related, does an equivalence match of creativity and OE require strict commitment to a systems (and embodied) theory of creativity and rejection of isolated accounts?
  • What is the trade-off between OE, CC, and controllability in a dynamical system? How are the first two concepts related to an observer’s ability to control a system’s dynamics? Can the relationship of either one of the first two concepts illuminate the role of the other?
@@
Each question could only be elaborated on briefly; we encourage researchers interested in investigating them further to contact the authors for further discussion and clarification.
@@
The dual pursuits of synthesizing creativity in CC and OE in ALife involve understanding the rich histories of each research agenda, as briefly summarized in this paper. Our goal has not been to move toward a complete coalescence of creativity and OE but, through intentional dialogue about both the similarities and differences in approaching each challenge, we hope to arrive at a clearer vision of how best to further our individual aims by understanding as tightly interconnected with those of CC and ALife as neighbouring communities. To this end, we suggest a few questions that, if attended, would likely help push both communities toward a fuller understanding of these respective fields:
  • Can we translate research facets and goals from CC to OE research, and vice versa? For instance, an interesting exercise might be to imagine “creativity as it might be”, as previously proposed by Guckelsberger et al. [2017].
@@
A long history of exchange between ALife and CC [Bown and McCormack, 2009; Saunders, 2012; Bown, 2012; McCormack, 2012; Guckelsberger et al., 2017; McCormack, 2019] bears witness of this endeavour’s inspirational and transformational potential. The very positive feedback and acceptance at one of the primary venues of ALife research highlight the willingness of the ALife community to take this initiative forward; separate calls by CC researchers to bridge communities [e.g. Cook and Colton, 2018] support that it will also fall on fertile CC ground.
@@
The engineering of CC (CC) and OE (OE) have been identified as the “final frontier” and “the last grand challenge” of AI research, respectively. We set out to (1) revisit previously highlighted and (2) collect further connections between creativity and OE as central research topics within CC and ALife. We identify opportunities for both communities to learn from each other’s research heritage, progress, and challenges, and consequently (3) advocate the advancement of OE and CC research through interdisciplinary collaborations.
To promote accessibility also from outside of these two fields, we provided short introductions and historical context for progress in both CC and ALife. Complementing our review of existing with fresh connections, we further support that creativity and OE, as conceptualised in these disciplines, are not generally identical. The new connections drawn point at nuanced but fundamental differences, and enable their detailed exploration in future work.
The focus of this paper has been conceptual, motivated by the struggle to define creativity and OE as one salient commonality. Also common to both endeavours is the ongoing debate about standardised measures; future work should revisit measurement proposals from both fields for joint advancement, focusing on e.g. attempts to measure creativity and OE objectively or subjectively, and sensitive to or agnostic w.r.t. a specific domain. Finally, we see an opportunity for future work to reflect closely on which methodology and meta-goals drive the study of creativity and OE to broaden our toolkits and scientific perspectives.
Crucially, this future work should not be done in isolation; we most importantly sound the call for further interdisciplinary discussions and collaboration between researchers interested in (computational) creativity and OE. It is crucial to establish a dialogue, since doing so might provide additional context, insight, and directions moving forward. We hope that such endeavors will deepen our understanding of both research challenges and further elucidate the requirements to reproduce them in artificial systems – including reducing ambiguity in conceptualisation and measurement.
@@ Acknowledgements
We thank our anonymous reviewers for their excellent feedback and support for this paper’s mission. LS, AA, and OW were supported by Cross Compass, Ltd.